An exciting area of astrophysical research is the discovery of exoplanets, also known as extrasolar planets, which are planetHello I need help with this question pls
Answer #1

Gravitational constant = G = 6.67 x 10-11 N.m2/kg2

Mass of the star = M = 2.79 x 1030 kg

Mass of the exoplanet Draugr = m = 1.2 x 1023 kg

Radius of the orbit of the exoplanet Draugr = R = 2.85 x 1010 m

Radius of the exoplanet Draugr = Rp = 2.05 x 106 m

Orbital speed of the exoplanet Draugr = V

The centripetal force for the circular motion of the exoplanet Draugr is provided by the gravitational force of the star on it.

mV GMm R = R2

12 GM/

GM V = VR

(6.67 + 10-11)(2.79 + 1030) 2.85 + 1010

V = 80805.81 m/s

Orbital period of the exoplanet Draugr = T

VT = 2\piR

(80805.81)T = 2\pi(2.85x1010)

T = 2.216 x 106 sec

Converting to days,

2.216 * 106 24 x 60 x 60

T = 25.6 days

Magnitude of acceleration of the exoplanet Draugr as it orbits the star = a

a = \frac{V^{2}}{R}

a = \frac{(80805.81)^{2}}{2.85\ast 10^{10}}

a = 0.229 m/s2

Free fall acceleration near the surface of the exoplanet Draugr = g

Gm 9= RM

g= (6.67 * 10-11) (1.2 * 1023 (2.05 * 106)2

g = 1.90 m/s2

14) Period of the motion of the exoplanet Draugr as it orbits the star = 25.6 days

15) Magnitude of acceleration of the exoplanet Draugr as it orbits the star = 0.229 m/s2

16) Free fall acceleration near the surface of the exoplanet Draugr = 1.90 m/s2